# From Similarity to Superiority: Channel Clustering for Time Series Forecasting

### v1

@@ Abstract
Time series forecasting has attracted significant attention in recent decades. Previous studies have demonstrated that the Channel-Independent (CI) strategy improves forecasting performance by treating different channels individually, while it leads to poor generalization on unseen instances and ignores potentially necessary interactions between channels. Conversely, the Channel-Dependent (CD) strategy mixes all channels with even irrelevant and indiscriminate information, which, however, results in oversmoothing issues and limits forecasting accuracy. There is a lack of channel strategy that effectively balances individual channel treatment for improved forecasting performance without overlooking essential interactions between channels. Motivated by our observation of a correlation between the time series model’s performance boost against channel mixing and the intrinsic similarity on a pair of channels, we developed a novel and adaptable Channel Clustering Module (CCM). CCM dynamically groups channels characterized by intrinsic similarities and leverages cluster identity instead of channel identity, combining the best of CD and CI worlds. Extensive experiments on real-world datasets demonstrate that CCM can (1) boost the performance of CI and CD models by an average margin of $2.4\%$ and $7.2\%$ on long-term and short-term forecasting, respectively; (2) enable zero-shot forecasting with mainstream time series forecasting models; (3) uncover intrinsic time series patterns among channels and improve interpretability of complex time series models.
Machine Learning, ICML
[…]
@@ 1 Introduction
Figure 1: The pipeline of applying Channel Clustering Module (CCM) to general time series models. (a) illustrates the general framework of most time series models. (b) illustrates two core modules modified when applying CCM to base models: cluster assigner and cluster-aware Feed Forward. The cluster assigner learns channel clustering based on intrinsic similarities and creates prototype embeddings for each cluster via a cross-attention mechanism. The clustering probabilities $\{p_{i,k}\}$ are subsequently used in the cluster-aware Feed Forward to average $\{\mathbf{\theta}_{k}\}_{k=1}^{K}$, which are layer weights assigned to $K$ clusters, obtaining weights $\mathbf{\theta}^{i}$ for the $i$-th channel. The learned prototypes retain pre-trained knowledge, enabling zero-shot forecasting on unseen samples in both univariate and multivariate scenarios.
@@
Time series forecasting has been attracting a surge of interest across diverse fields, ranging from economics, energy (Alvarez et al., 2010; Koprinska et al., 2018), weather (Angryk et al., 2020; Lam et al., 2023), to transportation planning (Li et al., 2015; Yin & Shang, 2016). The complexity of the task is heightened by factors including seasonality, trend, noise in the data, and potential cross-channel information.
@@
Despite the numerous deep learning time series models proposed recently (Chen et al., 2023; Zeng et al., 2022; Wen et al., 2022; Zhou et al., 2022b, 2021; Wu et al., 2021, 2022; Liu et al., 2022a), an unresolved challenge persists in the effective management of channel interaction within the forecasting framework (Li et al., 2023a; Peiwen & Changsheng, 2023). Previous works have explored two primary channel strategies: Channel-Independent (CI) and Channel-Dependent (CD) strategies. Channel-Independent (CI) strategy has shown promise in better forecasting performance by having individual models for each channel. However, a critical drawback is its limited generalizability and robustness on unseen channels (Han et al., 2023). Besides, it tends to overlook potential interactions between various channels. Conversely, Channel-Dependent (CD) strategy models all channels as a whole and captures intricate channel relations, while they tend to show oversmoothing and have trouble fitting to individual channels, especially when the similarity between channels is very low. Moreover, existing models typically treat univariate data in a CI manner, neglecting the interconnections between time series samples, even though these dependencies are commonly observed in real-world scenarios, such as stock market or weather forecasting (Ni & Xu, 2023; Yin et al., 2021; Jang & Matteson, 2023).
@@
Proposed work. To address the aforementioned challenges, we propose a Channel Clustering Module (CCM) that balances individual channel treatment and captures necessary cross-channel dependencies simultaneously. CCM is motivated by the following key observations: All models, CI and CD, heavily rely on channel identity information. The level of reliance is anti-correlated with the similarity between channels (see Sec. 4.1 for an analysis). This intriguing phenomenon alludes to the model’s analogous behavior on similar channels. The proposed CCM thereby involves the strategic clustering of channels into cohesive clusters, where intra-cluster channels exhibit a higher degree of similarity, effectively replacing channel identity with cluster identity. To capture the underlying time series patterns within these clusters, we employ cluster-aware Feed Forward to ensure individual management of each cluster. Moreover, CCM learns expressive prototype embeddings in the training phase, which enable zero-shot forecasting on unseen samples by grouping them into appropriate clusters.
@@
CCM is a plug-and-play solution that is adaptable to most mainstream time series models. We evaluate the effectiveness of CCM on four different time series backbones (aka. base models), TSMixer (Chen et al., 2023), DLinear (Zeng et al., 2022), PatchTST (Nie et al., 2022), and TimesNet (Wu et al., 2022). Extensive experiments verify the superiority of CCM in long-term and short-term forecasting benchmarks, achieving an average margin of $2.4\%$ and $7.2\%$, respectively. Additionally, we collect stock data from a diverse range of companies to construct a new stock univariate dataset. Leveraging information from intra-cluster samples, CCM consistently shows a stronger ability to accurately forecast stock prices within the dynamic and intricate stock market. Moreover, CCM enhances zero-shot forecasting capacities of time series backbones in cross-domain scenarios, which further highlights the robustness and versatility of CCM.
@@
The contributions of this paper are: (1) We propose a novel and unified channel strategy, i.e., CCM, which is adaptable to most mainstream time series models. CCM explores the optimal trade-off between channel individual treatment and cross-channel modeling, (2) Through learning prototypes from clusters, CCM enables zero-shot forecasting on unseen samples in both univariate and multivariate scenarios, and (3) CCM demonstrates superiority in improving performance on long-term and short-term forecasting.
@@ 2.1 Time Series Forecasting Models
Traditional machine learning methods such as Prophet (Taylor & Letham, 2018; Triebe et al., 2021), ARIMA (Zhang, 2003) capture the trend component and seasonality in time series (Ahmed et al., 2010). As data availability continues to grow, deep learning methods revolutionized this field, introducing more complex and efficient models (Torres et al., 2021; Lim & Zohren, 2021). Convolutional Neural Networks (CNNs) (Wu et al., 2022; Liu et al., 2022a; Bai et al., 2018; Wan et al., 2019; Sen et al., 2019), have been widely adopted to capture local temporal dependencies. Recurrent Neural Networks (RNNs) (Hewamalage et al., 2021; Rangapuram et al., 2018; Smyl, 2020; Salinas et al., 2020; Bai et al., 2018) excel in capturing sequential information, yet they often struggle with longer sequences. Transformer-based models (Zhou et al., 2021; Liu et al., 2022b; Wu et al., 2021; Nie et al., 2022; Zhang & Yan, 2022; Liu et al., 2023b; Wen et al., 2022; Tang & Matteson, 2021; Zhou et al., 2022b; Liu et al., 2021; Feng et al., 2024), typically equipped with self-attention mechanisms (Vaswani et al., 2017), demonstrate their proficiency in handling long-range dependencies, although they require substantial computational resources. Recently, linear models (Oreshkin et al., 2019; Zhang et al., 2022; Das et al., 2023), e.g., DLinear (Zeng et al., 2022), TSMixer (Chen et al., 2023), have gained popularity for their simplicity and effectiveness in long-term time series forecasting, but they may underperform with non-linear and complex patterns. Besides, traditional tricks, including trend-seasonal decomposition (Zeng et al., 2022; Cleveland et al., 1990; Wen et al., 2020) and multi-periodicity analysis (Benaouda et al., 2006; Percival & Walden, 2000; Wu et al., 2022; Wang et al., 2022; Zhang et al., 2024; Yi et al., 2023; Zhou et al., 2022a) continue to play a crucial role in aiding in the preprocessing stage for advanced models.
@@ 2.2 Channel Strategies in Time Series Forecasting
Most deep learning models (Wu et al., 2021; Liu et al., 2021; Zhou et al., 2022b) adopt the Channel-Dependent (CD) strategy, aiming to harness the full spectrum of information across channels. Conversely, the Channel-Independent (CI) approaches (Nie et al., 2022; Zeng et al., 2022) build forecasting models for each channel independently. Prior works on CI and CD strategy (Han et al., 2023; Li et al., 2023a; Montero-Manso & Hyndman, 2021) present that CI has high capacity and low robustness, whereas CD is the opposite. Predicting residuals with regularization (PRReg) (Han et al., 2023) is thereby proposed to incorporate a regularization term in the objective to encourage smoothness in future forecasting. However, the essential challenge of channel strategies from the model design perspective has not been solved and it remains challenging to develop a balanced channel strategy with both high capacity and robustness.
Prior research has explored effective clustering of channels to improve the predictive capabilities in diverse applications, including image classification (Jiang et al., 2010), natural language processing (NLP) (Marin et al., 2023; George & Sumathy, 2023), anomaly detection (Li et al., 2012; Syarif et al., 2012; Gunupudi et al., 2017). For instance, in traffic prediction (Ji et al., 2023; Liu et al., 2023a), clustering techniques have been proposed to group related traffic regions to capture intricate spatial patterns. Despite the considerable progress in these areas, the potential and effect of channel clustering in time series forecasting remain under-explored.
@@ 3 Preliminaries
Time Series Forecasting. Formally, let $X=[\boldsymbol{x}_{1},\ldots\boldsymbol{x}_{T}]\in\mathbb{R}^{T\times C}$ be a time series, where $T$ is the length of historical data. $\boldsymbol{x}_{t}\in\mathbb{R}^{C}$ represents the observation at time $t$. $C$ denotes the number of variates (i.e., channels). In the univariate case (i.e., $C=1$), the series is composed of single observations at each time step, whereas in the multivariate case (i.e., $C>1$), each $\boldsymbol{x}_{t}$ is a vector of multiple observations. The objective is to construct a predictive model $f$ that estimates the future values of the series, $Y=[\hat{\boldsymbol{x}}_{T+1},\ldots,\hat{\boldsymbol{x}}_{T+H}]\in\mathbb{R}^{H\times C}$, where $H$ is the forecasting horizon. We use $X_{[:,i]}\in\mathbb{R}^{T}$ ($X_{i}$ for simplicity) to denote the $i$-th channel in the time series.
Channel Dependent (CD) and Channel Independent (CI). The CI strategy models each channel $X_{i}$ separately and ignores the potential channel interactions. This approach is typically denoted as $f^{(i)}:\mathbb{R}^{T}\rightarrow\mathbb{R}^{H}$ for $i=1,\cdots,C$, where $f^{(i)}$ gets access to the historical data specific to the $i$-th channel. In contrast, the CD strategy models all the channels as a whole, using a function $f:\mathbb{R}^{T\times C}\rightarrow\mathbb{R}^{H\times C}$. This strategy is essential in scenarios where channels are not just parallel data streams but are interrelated, such as in financial markets or traffic flows.
@@ 4 Proposed Method
In this work, we propose Channel Clustering Module (CCM), a model-agnostic method that is adaptable to most mainstream models. The pipeline of applying CCM is visualized in Figure 1. General time series models, shown in Figure 1(a), typically consist of three core components (Li et al., 2023a, b): an optional normalization layer (e.g., RevIN (Kim et al., 2021), SAN (Liu et al., 2023c)), temporal modules including linear layers, transformer-based, or convolutional backbones, and a feed-forward layer that forecasts the future values. Motivated by the empirical observation discussed in Sec. 4.1, CCM presents with a cluster assigner preceding the temporal modules, followed by a cluster-aware Feed Forward (Sec. 4.2). The cluster assigner implements channel clustering based on intrinsic similarities and employs a cross-attention mechanism to generate prototypes for each cluster, which stores the knowledge from the training set and endows the model with zero-shot forecasting capacities.
@@ 4.1 Motivation for Channel Similarity
To motivate our similarity-based clustering method, we provide a toy experiment to verify the following two assumptions: (1) Existing forecasting methods heavily rely on channel identity information. (2) This reliance clearly anti-correlates with channel similarity: for channels with high similarity, channel identity information is less important. Together, these two assumptions motivate us to design an approach that provides cluster identity instead of channel identity, combining the best of both worlds: high capacity and generalizability.
Table 1: Averaged performance gain from a channel identity information ($\Delta\mathcal{L}(\%)$) and Pearson Correlation Coefficients (PCC) between $\{\Delta\mathcal{L}_{ij}\}_{i,j}$ and $\{\textsc{Sim}(X_{i},X_{j})\}_{i,j}$. The values are averaged across all test samples.
@@
We select four recent and popular time series models with different backbones. TSMixer (Chen et al., 2023) and DLinear are linear models. PatchTST (Nie et al., 2022) is a transformer-based model with a patching mechanism and TimesNet (Wu et al., 2022) is a convolutional network that captures multi-periodicity in data. Among these, TSMixer and TimesNet utilize a Channel-Dependent strategy while DLinear and PatchTST adopt the Channel-Independent design. We train a time series model across all channels and evaluate the channel-wise Mean Squared Error (MSE) loss on the test set. Then, we repeat training while randomly shuffling channels in each batch. Note that for both CD and CI models, this means channel identity information will be removed. We report the average performance gain across all channels based on the random shuffling experiments (denoted as $\Delta\mathcal{L}(\%)$) in Table 1. We see that all models rely on channel identity information to achieve better performance.
Next, we show that this performance gap anti-correlates with channel similarity. We define channel similarity based on radial basis function kernels (Zhang et al., 2004) as
$\textsc{Sim}(X_{i},X_{j})=\text{exp}(\frac{-\|X_{i}-X_{j}\|^{2}}{2\sigma^{2}}),\vspace{-0.1cm}$ (1)
where $\sigma$ is a scaling factor. Note that the similarity is computed on the standardized time series to avoid scaling differences. More details are discussed in Appendix A. The performance difference in MSE from the random shuffling experiment for channel $i$ is denoted as $\Delta\mathcal{L}_{i}$. We define $\Delta\mathcal{L}_{ij}:=|\Delta\mathcal{L}_{i}-\Delta\mathcal{L}_{j}|$ and calculate the Pearson Correlation Coefficients between $\{\Delta\mathcal{L}_{ij}\}_{i,j}$ and $\{\textsc{Sim}(X_{i},X_{j})\}_{i,j}$, as shown in Table 1, consistently verifying our assumption.

### v2

@@ Abstract
Time series forecasting has attracted significant attention in recent decades. Previous studies have demonstrated that the Channel-Independent (CI) strategy improves forecasting performance by treating different channels individually, while it leads to poor generalization on unseen instances and ignores potentially necessary interactions between channels. Conversely, the Channel-Dependent (CD) strategy mixes all channels with even irrelevant and indiscriminate information, which, however, results in oversmoothing issues and limits forecasting accuracy. There is a lack of channel strategy that effectively balances individual channel treatment for improved forecasting performance without overlooking essential interactions between channels. Motivated by our observation of a correlation between the time series model’s performance boost against channel mixing and the intrinsic similarity on a pair of channels, we developed a novel and adaptable Channel Clustering Module (CCM). CCM dynamically groups channels characterized by intrinsic similarities and leverages cluster information instead of individual channel identities, combining the best of CD and CI worlds. Extensive experiments on real-world datasets demonstrate that CCM can (1) boost the performance of CI and CD models by an average margin of $2.4\%$ and $7.2\%$ on long-term and short-term forecasting, respectively; (2) enable zero-shot forecasting with mainstream time series forecasting models; (3) uncover intrinsic time series patterns among channels and improve interpretability of complex time series models  [^1] The code is available at https://github.com/Graph-and-Geometric-Learning/TimeSeriesCCM.
@@ 1 Introduction
Time series forecasting has attracted a surge of interest across diverse fields, ranging from economics, energy [1, 2], weather [3, 4], to transportation planning [5, 6]. The complexity of the task is heightened by factors including seasonality, trend, noise in the data, and potential cross-channel information.
@@
Despite the numerous deep learning time series models proposed recently [7, 8, 9, 10, 11, 12, 13, 14], an unresolved challenge persists in the effective management of channel interaction within the forecasting framework [15, 16]. Previous works have explored two primary channel strategies: Channel-Independent (CI) and Channel-Dependent (CD) strategies. The Channel-Independent (CI) strategy has shown promise in better forecasting performance by having individual models for each channel. However, a critical drawback is its limited generalizability and robustness on unseen channels [17]. Besides, it tends to overlook potential interactions between various channels. Conversely, the Channel-Dependent (CD) strategy models all channels as a whole and captures intricate channel relations, while they tend to show oversmoothing and have trouble fitting to individual channels, especially when the similarity between channels is very low. Moreover, existing models typically treat univariate data in a CI manner, neglecting the interconnections between time series samples, even though these dependencies are commonly observed and beneficial in real-world scenarios, such as stock market or weather forecasting [18, 19, 20].
Figure 1: The pipeline of applying Channel Clustering Module (CCM) to general time series models. (a) is the general framework of most time series models. (b) illustrates two modified modules when applying CCM: Cluster Assigner and Cluster-aware Feed Forward. Cluster Assigner learns channel clustering based on intrinsic similarities and creates prototype embeddings for each cluster via a cross-attention mechanism. The clustering probabilities $\{p_{i,k}\}$ are subsequently used in Cluster-aware Feed Forward to average $\{\mathbf{\theta}_{k}\}_{k=1}^{K}$, which are layer weights assigned to $K$ clusters, obtaining weights $\mathbf{\theta}^{i}$ for the $i$-th channel. The learned prototypes retain pre-trained knowledge, enabling zero-shot forecasting on unseen samples in both univariate and multivariate scenarios.
@@
Proposed work. To address the aforementioned challenges, we propose a Channel Clustering Module (CCM) that balances individual channel treatment and captures necessary cross-channel dependencies simultaneously. CCM is motivated by the key observations that CI and CD models typically rely on channel identity information. The level of reliance is anti-correlated with the similarity between channels (see Sec. 4.1 for an analysis). This intriguing phenomenon alludes to the model’s analogous behavior on similar channels. The proposed CCM thereby involves the strategic clustering of channels into cohesive clusters, where intra-cluster channels exhibit a higher degree of similarity. To capture the underlying time series patterns within these clusters, we employ cluster-aware Feed Forward to assign independent weights to each cluster and replace individual channel treatment with individual cluster treatment. Moreover, CCM learns expressive prototype embeddings in training, which enables zero-shot forecasting on unseen samples by grouping them into appropriate clusters.
@@
CCM is a plug-and-play solution that is adaptable to most mainstream time series models. We evaluate the effectiveness of CCM on four different time series backbones (aka. base models): TSMixer [7], DLinear [8], PatchTST [21], and TimesNet [13]. It can also be applied to other state-of-the-art models for enhanced performance. Extensive experiments verify the superiority of CCM in long-term and short-term forecasting benchmarks, achieving an average margin of $2.4\%$ and $7.2\%$, respectively. Additionally, we collect stock data from a diverse range of companies to construct a new stock univariate dataset. Leveraging information from intra-cluster samples, CCM consistently shows a stronger ability to accurately forecast stock prices in the dynamic and intricate stock market. Moreover, CCM enhances zero-shot forecasting capacities of time series backbones in cross-domain scenarios, which further highlights the robustness and versatility of CCM.
@@
The contributions of this paper are: (1) We propose a novel and unified channel strategy, i.e., CCM, which is adaptable to most mainstream time series models. CCM explores the optimal trade-off between channel individual treatment and cross-channel modeling, (2) CCM demonstrates superiority in improving performance on long-term and short-term forecasting, and (3) through learning prototypes from clusters, CCM enables zero-shot forecasting on unseen samples in both univariate and multivariate scenarios.
@@ 2.1 Time Series Forecasting Models
Traditional machine learning methods such as Prophet [22, 23], ARIMA [24] capture the trend component and seasonality in time series [25]. As data availability continues to grow, deep learning methods revolutionized this field, introducing more complex and efficient models [26, 27]. Convolutional Neural Networks (CNNs) [13, 14, 28, 29, 30], have been widely adopted to capture local temporal dependencies. Recurrent Neural Networks (RNNs) [31, 32, 33, 34, 28] excel in capturing sequential information, yet they often struggle with longer sequences. Transformer-based models [11, 35, 12, 21, 36, 37, 9, 38, 10, 39, 40], typically equipped with self-attention mechanisms [41], demonstrate their proficiency in handling long-range dependencies, although they require substantial computational resources. Recently, linear models [42, 43, 44], e.g., DLinear [8], TSMixer [7], have gained popularity for their simplicity and effectiveness in long-term time series forecasting, but they may underperform with non-linear and complex patterns. Besides, traditional tricks, including trend-seasonal decomposition [8, 45, 46] and multi-periodicity analysis [47, 48, 13, 49, 50, 51, 52] continue to play a crucial role in aiding in the preprocessing stage for advanced models.
@@ 2.2 Channel Strategies in Time Series Forecasting
Most deep learning models [12, 39, 10] adopt the Channel-Dependent (CD) strategy, aiming to harness the full spectrum of information across channels. Conversely, the Channel-Independent (CI) approaches [21, 8] build forecasting models for each channel independently. Prior works on CI and CD strategy [17, 15, 53, 54, 16] present that CI leads to higher capacity and lower robustness, whereas CD is the opposite. Predicting residuals with regularization (PRReg) [17] is thereby proposed to incorporate a regularization term in the objective to encourage smoothness in future forecasting. However, the essential challenge from the model design perspective has not been solved and it remains challenging to develop a balanced channel strategy. Prior research has explored effective clustering of channels to improve the predictive capabilities in diverse applications, including image classification [55], natural language processing (NLP) [56, 57], anomaly detection [58, 59, 60]. For instance, in traffic prediction [61, 62], clustering techniques have been proposed to group relevant traffic regions to capture intricate spatial patterns. Despite the considerable progress in these areas, the potential and effect of channel clustering in time series forecasting remain under-explored.
@@ 3 Preliminaries
Time Series Forecasting. Formally, let $X=[\boldsymbol{x}_{1},\ldots\boldsymbol{x}_{T}]\in\mathbb{R}^{T\times C}$ be a time series, where $T$ is the length of historical data. $\boldsymbol{x}_{t}\in\mathbb{R}^{C}$ represents the observation at time $t$. $C$ denotes the number of variates (i.e., channels). The objective is to construct a predictive model $f$ that estimates the future values of the series, $Y=[\hat{\boldsymbol{x}}_{T+1},\ldots,\hat{\boldsymbol{x}}_{T+H}]\in\mathbb{R}^{H\times C}$, where $H$ is the forecasting horizon. We use $X_{[:,i]}\in\mathbb{R}^{T}$ ($X_{i}$ for simplicity) to denote the $i$-th channel in the time series.
Channel Dependent (CD) and Channel Independent (CI). The CI strategy models each channel $X_{i}$ separately and ignores any potential cross-channel interactions. This approach is typically denoted as $f^{(i)}:\mathbb{R}^{T}\rightarrow\mathbb{R}^{H}$ for $i=1,\cdots,C$, where $f^{(i)}$ is specifically dedicated to the $i$-th channel. Refer to Appendix A.2 for more details. In contrast, the CD strategy models all the channels as a whole with a function $f:\mathbb{R}^{T\times C}\rightarrow\mathbb{R}^{H\times C}$. This strategy is essential in scenarios where channels are not just parallel data streams but are interrelated, such as in financial markets or traffic flows.
@@ 4 Proposed Method
In this work, we propose a Channel Clustering Module (CCM), a model-agnostic method that is adaptable to most mainstream time series models. The pipeline of applying CCM is visualized in Figure 1. General time series models, shown in Figure 1(a), typically consist of three core components [15, 63]: an optional normalization layer (e.g., RevIN [64], SAN [65]), temporal modules including linear layers, transformer-based, or convolutional backbones, and a feed-forward layer that forecasts the future values. Motivated by the empirical observation discussed in Sec. 4.1, CCM presents with a cluster assigner preceding the temporal modules, followed by a cluster-aware Feed Forward (Sec. 4.2). The cluster assigner implements channel clustering based on intrinsic similarities and employs a cross-attention mechanism to generate prototypes for each cluster, which stores the knowledge from the training set and endows the model with zero-shot forecasting capacities.
@@ 4.1 Motivation for Channel Similarity
To motivate our similarity-based clustering method, we conduct the following toy experiment. We select four recent and popular time series models with different backbones. TSMixer [7] and DLinear are linear models. PatchTST [21] is a transformer-based model with a patching mechanism and TimesNet [13] is a convolutional network that captures multi-periodicity in data. Among these, TSMixer and TimesNet utilize a Channel-Dependent strategy while DLinear and PatchTST adopt the Channel-Independent design. We train a time series model across all channels and evaluate the channel-wise Mean Squared Error (MSE) loss on the test set. Then, we repeat training while randomly shuffling channels in each batch. Note that for both CD and CI models, this means channel identity information will be removed. We report the average performance gain in terms of MSE loss across all channels based on the random shuffling experiments (denoted as $\Delta\mathcal{L}(\%)$) in Table 1.
Table 1: Averaged performance gain from channel identity information ($\Delta\mathcal{L}(\%)$) and Pearson Correlation Coefficients (PCC) between $\{\Delta\mathcal{L}_{ij}\}_{i,j}$ and $\{\textsc{Sim}(X_{i},X_{j})\}_{i,j}$. The values are averaged across all test samples.
@@
We attribute the models’ performance decrease in the random shuffling experiments to the loss of channel identity information. We see that all models rely on channel identity information to achieve better performance. Next, we define channel similarity based on radial basis function kernels [66] as
$\textsc{Sim}(X_{i},X_{j})=\text{exp}(\frac{-\|X_{i}-X_{j}\|^{2}}{2\sigma^{2}}),$ (1)
where $\sigma$ is a scaling factor. Note that the similarity is computed on the standardized time series to avoid scaling differences. More details are discussed in Appendix A.1. The performance difference in MSE from the random shuffling experiment for channel $i$ is denoted as $\Delta\mathcal{L}_{i}$. We define $\Delta\mathcal{L}_{ij}:=|\Delta\mathcal{L}_{i}-\Delta\mathcal{L}_{j}|$ and calculate the Pearson Correlation Coefficients (PCC) between $\{\Delta\mathcal{L}_{ij}\}_{i,j}$ and $\{\textsc{Sim}(X_{i},X_{j})\}_{i,j}$, as shown in Table 1. The toy example verifies the following two assumptions: (1) Existing forecasting methods heavily rely on channel identity information. (2) This reliance clearly anti-correlates with channel similarity: for channels with high similarity, channel identity information is less important. Together, these two assumptions motivate us to design an approach that provides cluster identity instead of channel identity, combining the best of both worlds: high capacity and generalizability.